OpenStudy (anonymous):

Bari wants to know how she did in math class this month. These are Bari's quiz scores. 8, 17, 9, 3, 13

OpenStudy (anonymous):

are you trying to find the mean score?

OpenStudy (anonymous):

average=(8+17+9+3+13)/5=50/5=10